# Reliable Decision Support with LLMs: A Framework for Evaluating Consistency in Binary Text Classification Applications

### v1

@@ 4.1 Planning Phase
Our case study implements the psychometric approach outlined in Section 3.1.1 with a practical adjustment. We applied a Šidák correction to control the family-wise error rate when conducting multiple comparisons. We adjusted the base significance level $\alpha$ to $\alpha^{*}=(1-0.9^{1/7})\approx 0.0149$, where 0.90 represents our target confidence level, and 7 represents the maximum number of LLMs we would realistically explore in the intra-rater reliability scenario. This adjustment produces a more conservative critical value for $z_{\alpha^{*}/2}$ to compensate for the increased risk of type I error in multiple model comparisons.
We extracted inverse $C$ values from Gwet’s agreement tables [48] based on binary classification ($q=2$) and rater counts ($r=5,7,9$). We used odd values of $r$ to reduce the likelihood of ties. Figure 2 displays the resulting sample sizes across reliability metrics. Brennan-Prediger required the largest sample ($n_{final}=1{,}317$), Percent Agreement a moderate size (847), and Gwet’s AC1 the smallest (216). All metrics show reduced sample size needs as the number of raters increases, reflecting the statistical efficiency of additional raters.
@@
We selected 14 LLMs (depicted in Figure 3) for evaluation. Our selection spans seven prominent LLM providers: Anthropic, OpenAI, Google, Microsoft, DeepSeek, Meta, and Cohere. We evaluated two models per provider, allowing a direct comparison of performance and cost-effectiveness within each ecosystem. This design enabled us to assess whether higher-tier models offer measurable advantages over their smaller, less expensive counterparts—a trade-off between compute cost and classification performance.
@@
Our prompt design for the case study builds upon the manual CoT and few-shot learning approaches discussed in Section 2. Following standard practice, we developed a two-part prompt structure: (a) a system prompt with instructions and examples and (b) a user prompt that delivers the specific news article information. In the system prompt, we direct the LLM to first read the article’s title and text. Then, we instruct the LLM to analyze the impact of this information on the company and identify key factors that may influence investor sentiment. The LLM is further instructed to assess the overall tone and its potential impact on the stock price. We also present it with definitions of what we mean by “positive” and “negative” sentiment and an output format that we hope it follows to facilitate our parsing of information. Our system prompt is completed with two carefully constructed examples, where we demonstrate the expected reasoning process and output format. The user prompt template passes the news article’s title, full text, and stock ticker symbol to the LLM.
@@ 4.2 Data Collection Phase
We extracted stock news articles using the StockNewsAPI [64] to investigate LLM reliability in financial sentiment classification. Data was retrieved on March 14, 2025, at approximately 12:30 EDT. Our API query used multiple parameters to obtain a pertinent dataset. We requested articles from January 05, 2025, to March 13, 2025. We selected this recent time frame to ensure these articles were not included in any LLM training data, as all evaluated models have knowledge cut-off dates before 2025. We limited the results to tickers on the NYSE and NASDAQ exchanges and US-based results. This focus on major US exchanges ensures market liquidity, standardized reporting practices, and comprehensive news coverage. We queried only positive and negative sentiment articles to maximize separation between the dichotomous sentiment classes. We collected article-type content (i.e., no videos), set 100 items per page, and iterated through multiple pages to gather our dataset.
The initial dataset contained 10,000 articles with attributes: article URL, headline, full-text content, publishing source, publication date, related topics, API-provided sentiment classification, content type (all values = “article”), and associated stock ticker symbols. These 10,000 articles contained 7,088 positive and 2,912 negative samples. The API labeled these sentiment tags through keyword analysis of each article’s title and subtitle [64, per their FAQ page]. Despite this straightforward approach, these labels are widely used as reference values in the financial sentiment analysis literature. Hence, in our case study, we consider the API-provided sentiment labels as the “benchmark” for our evaluation.
We performed several preprocessing steps on the initial dataset. We removed articles with multiple tickers to retain only news focused on a single stock. This ensures sentiment classification applies to specific companies rather than mixing signals across multiple entities. We implemented a sampling strategy that prevents the same ticker from appearing multiple times within the same sentiment category. Thus, a ticker can appear at most twice in our preprocessed dataset. This was followed by random downsampling of the dataset to 1,350 articles, equally divided between positive and negative. Our final dataset for LLM-based classification contained 1,183 unique stock tickers.
@@
Our case study examined sentiment classification across $n=1,350$ financial news articles using a diverse set of 14 LLMs (from seven providers). Each model processed each article five times, yielding $6,750\ (1,350\times 5)$ prompt completions per LLM. To ensure consistency, we constructed prompts by combining the article title, full text, and ticker symbol with a standardized system message. The model suite encompassed commercial API-based models and locally hosted open-weight alternatives deployed via Ollama. We configured all runs with a set temperature of 0 to minimize the models’ creativity. Furthermore, we set a 3000-token limit to ensure all models have the same output configurations. We implemented a retry mechanism for up to three failed attempts per request in case of any API connection and rate-limiting issues. For sentiment extraction, we used regular-expression-based pattern-matching techniques to extract the binary classification labels (“positive” or “negative”) and developed specialized logic to resolve inconsistent response formatting.
@@
Figure 4: Comparison of LLM inference times and generated sentiment distributions.
@@ 4.3 Reliability Assessment
Most models demonstrated remarkably high intra-rater consistency. The NA-penalized agreement distributions showed that most models achieve 100% agreement on 88–98% of articles, with mean agreement rates exceeding 0.93 for all models. The other reliability metrics reinforce these trends. Models such as gemma3:27B and claude-3-5-haiku achieved values about 0.99 across Brennan-Prediger, Conger’s Kappa, and Krippendorff’s Alpha. Interestingly, intra-LLM consistency was similar across smaller and larger models within the same provider, suggesting diminishing returns from increased model size, at least with respect to self-consistency, except in the case of the deepseek models.
@@ 5 Discussion
Claude-3-5-haiku, demonstrated performance metrics comparable to larger, more expensive models such as GPT-4o and Claude-3-7-sonnet. This challenges the “fear of missing out” (FOMO) mentality that drives many organizations to continually upgrade to newer, larger models without clear evidence of improved task performance. This finding is particularly relevant for firms seeking to scale LLM solutions while managing computational and financial costs. The prevailing industry trend towards ever larger models has often overshadowed the potential of smaller, faster models to deliver high-quality results, especially for narrowly defined tasks like binary classification [7]. Our observations align with prior studies questioning the cost-effectiveness of larger models for routine classification tasks [14]. These results reinforce the importance of selecting models based on task-specific performance evaluations rather than assumptions based on model size.
@@
Claude-3-5-haiku—achieved high levels of accuracy against the benchmark model, their annotations did not align as strongly with real-world financial market movements (i.e., the external criterion). This discrepancy underscores that while LLMs can produce semantically consistent labels, these labels may not always reflect actual behavioral outcomes such as stock price changes. This finding echoes prior research (e.g., Aguda et al. [6]) suggesting that LLMs may excel at tasks requiring semantic alignment but struggle with tasks requiring complex causal inference or behavioral prediction. As such, LLM-generated annotations may be more appropriate for descriptive or classification-oriented applications than for predictive modeling, unless additional task-specific tuning is applied.
In sum, while careful experimental design, appropriate sample size planning, and nuanced interpretation of performance metrics remain critical, as emphasized by [2], researchers and practitioners must also carefully distinguish between tasks involving semantic judgment and those requiring true predictive inference.
@@ 6.1 Theoretical Implications
This study advances the theoretical foundations of LLM reliability evaluation by introducing a scientific framework for using LLMs as structured raters in classification tasks. First, it contributes to growing work on replicable LLM systems by proposing a statistical method for determining minimum sample sizes for intra-rater reliability, addressing a gap where prior studies relied on arbitrary choices (e.g., [65]). This psychometrically informed approach bridges empirical LLM evaluation with reproducible social science methods. It shifts the focus from the accuracy of the model [66] to the stability of the measurement, viewing LLMs as bounded probabilistic annotators. This perspective aligns the evaluation of LLMs with long-standing measurement theory practices that emphasize consistency, validity, and reproducibility. Second, the empirical findings challenge the assumption that larger models offer greater reliability. Smaller models (e.g., llama3.2:1B, claude-3-5-haiku) matched or exceeded larger ones in consistency and accuracy, suggesting model architecture, prompt adherence, and task specificity may matter more than scale. This invites refinement of LLM scaling theories, proposing that emergent abilities may not extend uniformly to structured classification tasks. Third, the disconnect between high consistency and low predictive validity highlights a theoretical boundary. While consistency is necessary for reliability, it is not sufficient. Highly consistent models like Gemma 3 and Claude 3 failed to predict stock movements, supporting prior claims about LLMs’ domain-specific limitations [10, 13]. This underscores the need to distinguish tasks requiring semantic consistency from those requiring real-world prediction, avoiding overgeneralization across fundamentally different task types. Finally, the framework’s modular design supports its application across domains such as legal text labeling, medical triage, and educational content tagging. As LLMs gain traction in industry, structured evaluations are critical for user trust, consistent with findings from human-AI collaboration research [17].
@@ 6.2 Practical Implications
This study offers actionable guidance for practitioners applying LLMs to classification and annotation tasks. First, the proposed framework offers a replicable, resource-efficient method for validating LLMs before full deployment. By adapting psychometric reliability analyses, practitioners can establish empirically justified consistency thresholds and benchmark models without relying on subjective evaluations. Second, smaller, less expensive LLMs proved highly viable, achieving reliability and accuracy comparable to larger models. With small models offering 5–29 times cost reductions [67], organizations in resource-constrained environments can prioritize task-specific suitability over defaulting to flagship or high-parameter models. Third, the framework emphasizes the need to handle invalid responses explicitly—using NA-penalized and NA-dropped strategies—to ensure reliability assessments reflect real-world conditions. Practitioners can apply these approaches to build more robust, production-ready annotation pipelines. Fourth, while LLM annotations reliably replicate benchmark labels, their failure to predict future outcomes highlights a key limitation. LLMs are well-suited for structured classification tasks, but should not be assumed to provide predictive insights into stochastic real-world phenomena without additional modeling layers or complementary methods. Finally, the open-source code and replicable experimental setup enable rapid adoption and adaptation by research groups and industry teams, promoting best practices in evaluating LLM-based systems for annotation and decision support. Together, these insights offer a practical roadmap for organizations seeking to deploy LLMs responsibly, efficiently, and with measurable reliability.
@@ 7 Limitations and Future Research
While this study presents a rigorous framework for evaluating LLM consistency, several limitations suggest directions for future research. First, our analysis focused on binary classification, whereas many real-world tasks involve multiclass settings. Extending the framework to multiclass contexts introduces challenges, such as adapting reliability metrics and recalibrating sample size estimates for increased label complexity. Second, generalizability is limited by the case study’s focus on financial sentiment. Applying the framework across domains will clarify whether observed reliability patterns hold more broadly and how model size influences performance in varied tasks. Third, handling non-compliant LLM responses remains a challenge. We used a lenient regular-expression approach to extract misformatted labels; alternatives include penalizing non-conformity or constraining outputs to formats like JSON [68]. Finally, the framework currently applies only to text-based tasks. Future work should explore its use with non-text modalities, such as image classification via vision transformers, to assess whether reliability patterns extend to multimodal systems—particularly relevant for Industry 4.0 and 5.0 applications [69].
@@ 8 Conclusion
LLMs are increasingly used for text annotation and classification in decision support systems, where rigorous evaluation is critical. This study presents a four-phase framework for assessing LLM consistency, reliability and validity in binary classification, grounded in psychometric principles for model selection, sample sizing, and reliability analysis. In a financial sentiment case study involving 14 LLMs and 1,350 articles, both small and large models achieved high intra- and inter-rater reliability—with smaller models often matching or exceeding larger ones in consistency and cost-efficiency. This finding can help organizations optimize resources. However, while LLMs replicated benchmark sentiment labels, they failed to predict stock movements, underscoring the gap between annotation reliability and real-world predictive validity. LLMs are well-suited for structured annotation but require careful evaluation in predictive contexts. The proposed framework offers replicable tools for deploying LLMs in scalable, trustworthy annotation workflows.
@@
Our knitted Markdown, available at https://fmegahed.github.io/research/llm_consistency/llm_consistency.html, integrates the and Python code to evaluate LLM consistency in binary text annotation tasks. It includes the code and the corresponding results, such as figures and tables, providing a self-contained resource for implementing our framework. The accompanying GitHub repository, https://github.com/fmegahed/llm_consistency, contains the stored input files and the outputs obtained from our analysis, ensuring transparency and accessibility.

### v2

@@ 4.1 Planning Phase
Our case study implements the psychometric approach outlined in Section 3.1.1 with a practical adjustment. We applied a Šidák correction to control the family-wise error rate when conducting multiple comparisons. We adjusted the base significance level $\alpha$ to $\alpha^{*}=(1-0.9^{1/7})\approx 0.0149$, where 0.90 represents our target confidence level, and 7 represents the maximum number of LLMs we would realistically explore in the intra-rater reliability scenario. We chose seven as a conservative upper bound on the number of LLMs likely to pass the intra-rater reliability step. This ensures that the adjusted significance level reflects the maximum set of pairwise comparisons that practitioners would reasonably conduct in this phase. This adjustment produces a more conservative critical value for $z_{\alpha^{*}/2}$ to compensate for the increased risk of type I error in multiple model comparisons.
We extracted inverse $C$ values from Gwet’s agreement tables [gwet2021handbook] based on binary classification ($q=2$) and rater counts ($r=5,7,9$). We used odd values of $r$ to reduce the likelihood of ties. Figure 2 displays the resulting sample sizes across reliability metrics. Brennan-Prediger required the largest sample ($n_{final}=1{,}317$), Percent Agreement a moderate size (847), and Gwet’s AC1 the smallest (216). All metrics show reduced sample size needs as the number of raters increases, reflecting the statistical efficiency of additional raters.
@@
We selected 14 LLMs (depicted in Figure 3) for evaluation. Our selection spans seven prominent LLM providers: Anthropic, OpenAI, Google, Microsoft, DeepSeek, Meta, and Cohere. We evaluated two models per provider, allowing a direct comparison of performance and cost-effectiveness within each ecosystem. This design enabled us to assess whether higher-tier models offer measurable advantages over their smaller, less expensive counterparts (i.e., is there a trade-off between compute cost and classification performance).
@@
Our prompt design for the case study builds upon the manual CoT and few-shot learning approaches discussed in Section 2. Following standard practice, we developed a two-part prompt structure: (a) a system prompt with instructions and examples and (b) a user prompt that delivers the specific news article information. In the system prompt, we direct the LLM to first read the article’s title and text. Then, we instruct the LLM to analyze the impact of this information on the company and identify key factors that may influence investor sentiment. The LLM is further instructed to assess the overall tone and its potential impact on the stock price. We also present it with definitions of what we mean by “positive” and “negative” sentiment and an output format that we hope it follows to facilitate our parsing of information. Our system prompt is completed with two carefully constructed examples, where we demonstrate the expected reasoning process and output format. The user prompt template passes the news article’s title, full text, and stock ticker symbol to the LLM. To ensure fair comparisons, the prompt structure, including the system instructions, worked examples, and user template, was kept fixed across all 14 LLMs, and no model- or API-specific adaptations were required.
@@ 4.2 Data Collection Phase
We extracted stock news articles using the StockNewsAPI [StockNewsAPI] to investigate LLM reliability in financial sentiment classification. Data was retrieved on March 14, 2025, at approximately 12:30 EDT. Our API query used multiple parameters to obtain a pertinent dataset. We requested articles from January 05, 2025, to March 13, 2025. We selected this recent time frame to ensure these articles were not included in any LLM training data, as all evaluated models have knowledge cut-off dates before 2025. We limited the results to tickers on the NYSE and NASDAQ exchanges and US-based results. This focus on major US exchanges ensures market liquidity, standardized reporting practices, and comprehensive news coverage. We queried only positive and negative sentiment articles to maximize separation between the dichotomous sentiment classes. We collected article-type content (i.e., no videos), set 100 items per page, and iterated through multiple pages to gather our dataset.
The initial dataset contained 10,000 articles with attributes: article URL, headline, full-text content, publishing source, publication date, related topics, API-provided sentiment classification, content type (all values = “article”), and associated stock ticker symbols. These 10,000 articles contained 7,088 positive and 2,912 negative samples. The API labeled these sentiment tags through keyword analysis of each article’s title and subtitle [StockNewsAPI, per their FAQ page]. Despite this straightforward approach, these labels are widely used as reference values in the financial sentiment analysis literature. Hence, in our case study, we consider the API-provided sentiment labels as the “benchmark” for our evaluation.
We performed several preprocessing steps on the initial dataset. First, we removed articles associated with multiple tickers to ensure that each article corresponds to a single company. Next, within each sentiment category (positive and negative), we enforced ticker-level uniqueness by retaining at most one article per ticker. This prevents multiple occurrences of the same ticker within a sentiment class and avoids overweighting companies with higher news volume. Under this design, a ticker may appear once in the positive set and once in the negative set, yielding a maximum of two occurrences across the preprocessed dataset. This was followed by random downsampling of the dataset to 1,350 articles, equally divided between positive and negative. Our final dataset for LLM-based classification contained 1,183 unique stock tickers.
@@
Our case study examined sentiment classification across $n=1,350$ financial news articles using a diverse set of 14 LLMs (from seven providers). Each model processed each article five times, yielding $6,750\ (1,350\times 5)$ prompt completions per LLM. To ensure consistency, we constructed prompts by combining the article title, full text, and ticker symbol with a standardized system message. The model suite encompassed commercial API-based models and locally hosted open-weight alternatives deployed via Ollama. We configured all runs with a set temperature of 0 to minimize the models’ creativity. However, our proposed framework is flexible and allows users to adjust the temperature according to their application needs, including experiments that assess how model consistency changes under different temperature settings. Furthermore, we set a 3000-token limit to ensure all models have the same output configurations. We implemented a retry mechanism for up to three failed attempts per request in case of any API connection and rate-limiting issues. For sentiment extraction, we used regular-expression-based pattern-matching techniques to extract the binary classification labels (“positive” or “negative”) and developed specialized logic to resolve inconsistent response formatting.
@@
Figure 4: Comparison of LLM inference times and generated sentiment distributions. An “invalid” label is assigned when the model output cannot be mapped to the required “positive” or “negative” categories. For example, the model responds with terms like “neutral” or “unsure,” or when the output cannot be parsed into either label.
@@ 4.3 Reliability Assessment
Most models demonstrated remarkably high intra-rater consistency. The NA-penalized agreement distributions showed that most models achieve 100% agreement on 88–98% of articles, with mean agreement rates exceeding 0.93 for all models. The other reliability metrics reinforce these trends. Models such as gemma3:27B and claude-3-5-haiku achieved values about 0.99 across Brennan-Prediger, Conger’s Kappa, and Krippendorff’s Alpha. Interestingly, for our case study, the intra-LLM consistency was similar across smaller and larger models within the same provider. This suggests diminishing returns from increased model size, at least with respect to self-consistency, except in the case of the deepseek models.
@@ 5 Discussion
Claude-3-5-haiku, demonstrated performance metrics comparable to larger, more expensive models such as GPT-4o and Claude-3-7-sonnet for our financial news sentiment annotation task. This challenges the “fear of missing out” (FOMO) mentality that drives many organizations to continually upgrade to newer, larger models without clear evidence of improved task performance. This finding is particularly relevant for firms seeking to scale LLM solutions while managing computational and financial costs. The prevailing industry trend towards ever larger models has often overshadowed the potential of smaller, faster models to deliver high-quality results, especially for narrowly defined tasks like binary classification [chiang2023can]. Our observations align with prior studies questioning the cost-effectiveness of larger models for routine classification tasks [reiss2023testing]. These results reinforce the importance of selecting models based on task-specific performance evaluations rather than assumptions based on model size.
@@
Claude-3-5-haiku—achieved high levels of accuracy against the benchmark model, their annotations did not align as strongly with real-world financial market movements (i.e., the external criterion). This discrepancy underscores that while LLMs can produce semantically consistent labels, these labels may not always reflect actual behavioral outcomes such as stock price changes. This finding echoes prior research (e.g., Aguda et al. [aguda2024large]) suggesting that LLMs may excel at tasks requiring semantic alignment but struggle with tasks requiring complex causal inference or behavioral prediction. As such, LLM-generated annotations may be more appropriate for descriptive or classification-oriented applications than for predictive modeling, unless additional task-specific tuning is applied.
It is important to note that both validity references used in our case study are imperfect ground truths. The API labels, while commonly used in financial sentiment research, are generated using rule-based keyword heuristics and may not fully capture nuanced investor sentiment. Likewise, short-horizon market returns reflect a complex mixture of information, noise, and macroeconomic dynamics, which limits their ability to serve as a definitive behavioural criterion. These limitations do not undermine the proposed framework; rather, they illustrate a central challenge in LLM validation studies, where reference labels and external criteria are themselves subject to measurement error. Our framework is designed to accommodate alternative or high-quality validity references in future applications, such as expert-annotated datasets or multi-day event study outcomes. Our case study demonstrates how validity assessment can be carried out even when ground truths are imperfect.
In sum, while careful experimental design, appropriate sample size planning, and nuanced interpretation of performance metrics remain critical, as emphasized by [gilardi2023chatgpt], researchers and practitioners must also carefully distinguish between tasks involving semantic judgment and those requiring true predictive inference.
@@ 6.1 Theoretical Implications
This study advances the theoretical foundations of LLM reliability evaluation by introducing a scientific framework for using LLMs as structured raters in classification tasks. First, it contributes to growing work on replicable LLM systems by proposing a statistical method for determining minimum sample sizes for intra-rater reliability, addressing a gap where prior studies relied on arbitrary choices (e.g., [zhou2021breaking]). This psychometrically informed approach bridges empirical LLM evaluation with reproducible social science methods. It shifts the focus from the accuracy of the model [kazari2025scaling] to the stability of the measurement, viewing LLMs as bounded probabilistic annotators. This perspective aligns the evaluation of LLMs with long-standing measurement theory practices that emphasize consistency, validity, and reproducibility. Second, the empirical findings challenge the assumption that larger models offer greater reliability. Smaller models (e.g., llama3.2:1B, claude-3-5-haiku) matched or exceeded larger ones in consistency and accuracy, suggesting model architecture, prompt adherence, and task specificity may matter more than scale. This invites refinement of LLM scaling theories, proposing that emergent abilities may not extend uniformly to structured classification tasks. Third, the disconnect between high consistency and low predictive validity highlights a theoretical boundary. While consistency is necessary for reliability, it is not sufficient. Highly consistent models like Gemma 3 and Claude 3 failed to predict stock movements, supporting prior claims about LLMs’ domain-specific limitations [tornberg2023chatgpt, stureborg2024large]. This underscores the need to distinguish tasks requiring semantic consistency from those requiring real-world prediction, avoiding overgeneralization across fundamentally different task types. Finally, the framework’s modular design supports its application across domains such as legal text labeling, medical triage, and educational content tagging. As LLMs gain traction in industry, structured evaluations are critical for user trust, consistent with findings from human-AI collaboration research [liu2024poliprompt].
@@ 6.2 Practical Implications
This study offers actionable guidance for practitioners applying LLMs to classification and annotation tasks. First, the proposed framework offers a replicable, resource-efficient method for validating LLMs before full deployment. By adapting psychometric reliability analyses, practitioners can establish empirically justified consistency thresholds and benchmark models without relying on subjective evaluations. Second, smaller, less expensive LLMs proved highly viable, achieving reliability and accuracy comparable to larger models. With small models offering 5–29 times cost reductions [irugalbandara2024scaling], organizations in resource-constrained environments can prioritize task-specific suitability over defaulting to flagship or high-parameter models. Third, the framework emphasizes the need to handle invalid responses explicitly—using NA-penalized and NA-dropped strategies—to ensure reliability assessments reflect real-world conditions. Practitioners can apply these approaches to build more robust, production-ready annotation pipelines. Fourth, while LLM annotations reliably replicate benchmark labels, their failure to predict future outcomes highlights a key limitation. LLMs are well-suited for structured classification tasks, but should not be assumed to provide predictive insights into stochastic real-world phenomena without additional modeling layers or complementary methods. Finally, the open-source code and replicable experimental setup enable rapid adoption and adaptation by research groups and industry teams, promoting best practices in evaluating LLM-based systems for annotation and decision support. Together, these insights offer a practical roadmap for organizations seeking to deploy LLMs responsibly, efficiently, and with measurable reliability.
@@ 7 Limitations and Future Research
While this study presents a rigorous framework for evaluating LLM consistency, several limitations suggest directions for future research. First, our analysis focused on binary classification, whereas many real-world tasks involve multiclass settings. Extending the framework to multiclass contexts introduces challenges, such as adapting reliability metrics and recalibrating sample size estimates for increased label complexity. An additional direction for future research involves extending the framework to fuzzy or graded annotation settings, where LLMs provide confidence scores or degrees of membership rather than discrete class labels. Generalizability is limited by the case study’s focus on financial sentiment. Applying the framework across domains will clarify whether observed reliability patterns hold more broadly and how model size influences performance in varied tasks. Third, handling non-compliant LLM responses remains a challenge. We used a lenient regular-expression approach to extract misformatted labels; alternatives include penalizing non-conformity or constraining outputs to formats like JSON [guzman2024introduction]. Finally, the framework currently applies only to text-based tasks. Future work should explore its use with non-text modalities, such as image classification via vision transformers, to assess whether reliability patterns extend to multimodal systems—particularly relevant for Industry 4.0 and 5.0 applications [megahed2025adapting].
@@ 8 Conclusion
LLMs are increasingly used for text annotation and classification in decision support systems, where rigorous evaluation is critical. This study presents a four-phase framework for assessing LLM consistency, reliability, and validity in binary classification, grounded in psychometric principles for model selection, sample sizing, and reliability analysis. In a financial sentiment case study involving 14 LLMs and 1,350 articles, both small and large models achieved high intra- and inter-rater reliability, with smaller models often matching or exceeding larger ones in consistency and cost-efficiency. This finding can help organizations optimize resources. However, while LLMs replicated benchmark sentiment labels, they failed to predict stock movements, underscoring the gap between annotation reliability and real-world predictive validity. LLMs are well-suited for structured annotation but require careful evaluation in predictive contexts. The proposed framework offers replicable tools for deploying LLMs in scalable, trustworthy annotation workflows.
@@
Our knitted Markdown, available at https://fmegahed.github.io/research/llm_consistency/llm_consistency.html, integrates the and Python code to evaluate LLM consistency in binary text annotation tasks. It includes the code and the corresponding results, such as figures and tables, providing a self-contained resource for implementing our framework.
To facilitate reproducibility and encourage further research in this area, we provide comprehensive materials in our GitHub repository https://github.com/fmegahed/llm_consistency. The repository contains four main components:
  1. (1)
    a data folder with the stock new articles and their labels.
  2. (2)
    a rmarkdown folder containing our   Markdown.
  3. (3)
    a figure folder storing code to compile the figures and results in this paper; and
  4. (4)
    a results folder storing all experimental outputs.